### v1

# On uniqueness of radial potentials for given Dirichlet spectra with distinct angular momenta

Abstract
We consider an inverse spectral problem for radial Schrödinger operators with singular potentials. First, we show that the knowledge of the Dirichlet spectra for infinitely many angular momenta $\ell$ satisfying a Müntz-type condition uniquely determines the potential. Then, in a neighborhood of the zero potential, we prove that the potential is uniquely determined by two Dirichlet spectra associated with distinct angular momenta in the cases $(\ell_{1},\ell_{2})=(0,1)$ and $(0,2)$. Our approach relies on an explicit analysis of the corresponding singular differential equation, combined with the classical Kneser–Sommerfeld formula. These results confirm, in the linearized setting and in these configurations, a conjecture originally formulated by Rundell and Sacks (2001).
1 Introduction
Inverse spectral problems for Schrödinger operators play a central role in quantum mechanics and mathematical physics, particularly in determining unknown potentials from spectral measurements. A classical example arises when the Schrödinger equation is considered in three spatial dimensions with a spherically symmetric potential in the unit ball of $\mathbb{R}^{3}$. In such cases, a standard separation of variables in spherical coordinates reduces the original partial differential equation to a family of one-dimensional singular Sturm-Liouville problems of the form
$\frac{d^{2}u}{dr^{2}}+\left(\frac{\ell(\ell+1)}{r^{2}}+q(r)\right)u=\lambda u,\qquad r\in(0,1),$ (1.1)
where $\ell\in\mathbb{N}$ is the angular momentum quantum number, and the radial potential $q$ is assumed to be real-valued and square-integrable on $(0,1)$. This equation is supplemented with a Dirichlet boundary condition at $r=1$ and a regularity condition at the origin, typically of the form $u(r)=O(r)$ as $r\to 0$, which ensures square-integrability of the solution and reflects the physical finiteness of the wavefunction near the origin.
Such reduced radial problems are not only mathematically rich but also physically relevant. They appear in a variety of contexts including quantum scattering in bounded domains, the study of acoustic modes in stellar interiors, and in zonal decompositions of Laplace-type operators on spheres. See, for instance, [5, 23].
A foundational result in regular inverse spectral theory is the Borg–Levinson theorem, which asserts that a potential can be uniquely recovered from the knowledge of the Dirichlet spectrum together with the associated norming constants, (that is, the Neumann data of the associated eigenfunctions; see below for further details). In the setting of singular Sturm–Liouville equations such as (1.1), Carlson [5] extended this result to include centrifugal singularities, establishing uniqueness under suitable conditions on the spectrum and norming data.
However, these classical results rely crucially on the availability of norming constants. In contrast, the present work investigates a different question, physically more relevant: can one determine the potential $q$ uniquely from the knowledge of the Dirichlet spectra corresponding to distinct values of $\ell$, without any norming constants?
Before addressing this question, let us recall some spectral properties of the underlying operator. For each fixed angular momentum $\ell\in\mathbb{N}$ and potential $q\in L^{2}(0,1)$, the Dirichlet spectrum is given by the zeros of an entire function (the so-called regular solution; see Appendix B), and consists of a countable sequence of simple, real eigenvalues $\{\lambda_{\ell,n}(q)\}_{n\geq 1}$ tending to infinity. These properties are well established in the literature [4, 5] and form the basis for both direct and inverse spectral analysis in the radial setting.
This observation naturally raises the question of whether combining spectral data from distinct angular momenta can lead to uniqueness. Indeed, the spectrum corresponding to a single value of $\ell$ does not determine the potential uniquely: the associated isospectral set is locally infinite-dimensional in $L^{2}(0,1)$ [5, 9, 19, 24]. By contrast, Carlson and Shubin [4] showed that combining spectral data from two distinct angular momenta $\ell_{1}\neq\ell_{2}$ significantly reduces the ambiguity. In particular, they proved that the corresponding isospectral set has finite local dimension when $\ell_{2}-\ell_{1}$ is odd.
Rundell and Sacks [23] went further and conjectured that the potential is uniquely determined by the Dirichlet spectra corresponding to any two distinct values of $\ell$, regardless of the parity of their difference. This reflects the fact that the angular dependence introduced via the centrifugal term $\ell(\ell+1)/r^{2}$ encodes genuinely distinct spectral information for each angular momentum channel, which can be leveraged to resolve the inverse problem even in the absence of norming constants.
Before turning to the more delicate case of two angular momenta, let us first consider the situation where Dirichlet spectra are known for infinitely many values of $\ell$. Our first theorem is a global uniqueness result and concerns the case where one knows the Dirichlet spectra for infinitely many angular momenta $\ell$ satisfying a Müntz-type condition. In fact, this result is a direct consequence of Theorem 2.1 established in [20] in the framework of scattering theory; see also ([6], Theorem 1.4) for local uniqueness results and [8] for an extension to magnetic fields. The proof of Theorem 1.1 is given in Appendix B, where we establish a correspondence between Dirichlet spectral data and scattering phases.
Theorem 1.1 (Global uniqueness).
Let $(\ell_{k})_{k\geq 1}$ be a strictly increasing sequence of positive integers such that
$\sum_{k=1}^{\infty}\frac{1}{\ell_{k}}=+\infty.$
Then the potential $q\in L^{2}(0,1)$ is uniquely determined by the Dirichlet spectra $\{\lambda_{\ell_{k},n}(q)\}_{k,n\geq 1}$.
Having established this global uniqueness result in the case of infinitely many spectra, we now turn to the more delicate situation involving only finitely many angular momenta. Because of its singular structure and the nonlinear dependence of the spectrum on the potential, the full inverse problem poses significant analytical challenges. A natural first step is to consider its linearization around a reference potential, typically $q\equiv 0$. Let $\psi_{\ell,n}(x)$ denote the $n$-th normalized eigenfunction associated with angular momentum $\ell$ and potential $q$. Then the Fréchet derivative of the corresponding eigenvalue with respect to $q$ is given by
$D_{q}\lambda_{\ell,n}(q)\,\zeta=\int_{0}^{1}\zeta(x)\,\psi_{\ell,n}(x)^{2}\,dx,$ (1.2)
where $\zeta$ denotes a perturbation of the potential, (see [23]). The norming constants are defined by $\gamma_{n}:=(\psi_{\ell,n}^{\prime}(1))^{2}$. In the linearized setting, the requirement that two potentials $q$ and $\check{q}$ yield the same eigenvalues for two angular momenta $\ell_{1}\neq\ell_{2}$ implies that $\zeta=\check{q}-q$ must be orthogonal to the family of squared eigenfunctions corresponding to both $\ell_{1}$ and $\ell_{2}$, evaluated at $q=0$.
In the unperturbed case $q\equiv 0$, the eigenfunctions of (1.1) can be expressed explicitly in terms of Bessel functions of the first kind. Specifically, for each $\ell\in\mathbb{N}$, the eigenfunctions take the form (up to normalization) [^1] For simplicity of exposition, we shall omit this inessential normalization throughout the present paper.
$\psi_{\ell,n}(r):=\sqrt{\frac{\pi r}{2}}\,J_{\nu}(j_{\nu,n}r),\qquad\nu=\ell+\tfrac{1}{2},\quad n\geq 1,$
where $J_{\nu}$ denotes the Bessel function of order $\nu$, and $j_{\nu,n}$ is its $n$-th positive zero. These functions satisfy Dirichlet boundary conditions at $r=1$, and regularity at the origin is ensured by the behavior $J_{\nu}(r)\sim r^{\nu}$ as $r\to 0$.
This leads to the problem of determining whether the span of the functions
$\left\{\psi_{\ell,n}^{2}(x)\,:\,n\geq 1,\ \ell\in\{\ell_{1},\ell_{2}\}\right\}$
is dense in $L^{2}(0,1)$. Establishing this completeness property in the linearized setting is a first step in order to prove the local uniqueness of the potential near $q=0$.
Even in the more favorable case where the angular momentum $\ell$ takes infinitely many values, it is not known whether the vector space spanned by the functions $\psi_{\ell,n}^{2}(x)$ is dense in $L^{2}(0,1)$, despite the fact that the problem is expected to be highly overdetermined. Although global uniqueness can now be derived by other methods, see Appendix B, the following result remains of independent interest.
Theorem 1.2 (Completeness).
Let $(\ell_{k})_{k\geq 1}$ be a strictly increasing sequence of positive integers such that
$\sum_{k=1}^{\infty}\frac{1}{\ell_{k}}=+\infty.$
Then the family
$\bigl\{\psi_{\ell_{k},n}^{2}(x):k\geq 1,\ n\geq 1\bigr\}$
is complete in $L^{2}(0,1)$.
Independently of the completeness result, the asymptotic behavior of the eigenvalues imposes further constraints on admissible perturbations. Specifically, as shown in [24, Proposition 3.1] (see also [23, Eq. (2.10)]), for each fixed angular momentum $\ell$, the following expansion holds uniformly on bounded subsets of $L^{2}(0,1)$:
$\lambda_{\ell,n}(q)=\left(n+\tfrac{\ell}{2}\right)^{2}\pi^{2}+\int_{0}^{1}q(x)\,dx-\ell(\ell+1)+\tilde{\lambda}_{\ell,n}(q),$ (1.3)
with $(\tilde{\lambda}_{\ell,n}(q)\in\ell_{\mathbb{R}}^{2}(\mathbb{N})$. Thus, if two potentials $q$ and $\check{q}$ produce identical eigenvalues for large $n$, their difference $\zeta=\check{q}-q$ must satisfy
$\int_{0}^{1}\zeta(x)\,dx=0.$
This condition reflects the fact that the mean of the potential appears as the subleading term in the asymptotic expansion and must therefore be preserved under any isospectral deformation.
We conclude this section with the main result of the paper. It shows that, in two specific cases, the Dirichlet spectra corresponding to two distinct angular momenta uniquely determine the potential in a neighborhood of the zero potential in the $L^{2}(0,1)$ topology.
Theorem 1.3 (Local uniqueness near the zero potential).
For $(\ell_{1},\ell_{2})=(0,1)$ or $(0,2)$, the knowledge of the Dirichlet spectra corresponding to angular momenta $\ell_{1}$ and $\ell_{2}$ uniquely determines the potential $q\in L^{2}(0,1)$ in a $L^{2}$-neighborhood of the zero potential.
Our approach is based on a detailed analysis of the spectral problem associated with (1.1) for varying values of $\ell$, in combination with the classical Kneser–Sommerfeld expansion. This representation plays a central role in establishing connections between the spectral data corresponding to different angular momenta, and in proving completeness results that support uniqueness in the linearized regime. We believe that this uniqueness result holds for any pair of distinct angular momenta $(\ell_{1},\ell_{2})$.
2 Proof of the main result
2.1 The spectral map
We define a spectral map involving the renormalized eigenvalues $\widetilde{\lambda}_{\ell,n}(q)$ defined by the asymptotic expansion (1.3) and associated with two distinct angular momenta. We then prove that this spectral map is an immersion at the zero potential and consequently locally injective at this point.
Let $\ell_{1}\neq\ell_{2}$ be two fixed non-negative integers. Consider the spectral map
$\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\longrightarrow\mathbb{R}\times\ell_{\mathbb{R}}^{2}(\mathbb{N})\times\ell_{\mathbb{R}}^{2}(\mathbb{N}),$
defined by
$\mathcal{S}_{\ell_{1},\ell_{2}}(q)=\left(\int_{0}^{1}q(x)\,dx,\;\big(\widetilde{\lambda}_{\ell_{1},n}(q)\big)_{n\geq 1},\;\big(\widetilde{\lambda}_{\ell_{2},n}(q)\big)_{n\geq 1}\right).$ (2.1)
The following proposition, proved in [24, Theorems 2.3 and 4.1] and based on (1.2), states that the spectral map is real-analytic and provides an explicit expression for its Fréchet differential at the zero potential. The subtraction of 1 in the integrals removes the constant part, already represented by the first coordinate of $\mathcal{S}_{\ell_{1},\ell_{2}}$.
Proposition 2.1.
Let $\ell_{1}\neq\ell_{2}$ be two fixed non-negative integers, and set $\nu_{k}=\ell_{k}+\tfrac{1}{2}$ for $k=1,2$. The spectral map
$\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\to\mathbb{R}\times\ell_{\mathbb{R}}^{2}(\mathbb{N})\times\ell_{\mathbb{R}}^{2}(\mathbb{N}),$
defined by
$\mathcal{S}_{\ell_{1},\ell_{2}}(q)=\left(\int_{0}^{1}q(r)\,dr,\;\big(\widetilde{\lambda}_{\ell_{1},n}(q)\big)_{n\geq 1},\;\big(\widetilde{\lambda}_{\ell_{2},n}(q)\big)_{n\geq 1}\right),$
is real-analytic. Moreover, for every $\zeta\in L^{2}(0,1)$, its Fréchet differential at the zero potential is given by
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}(\zeta)=\left(\int_{0}^{1}\zeta(r)\,dr,\quad\left(\int_{0}^{1}\zeta(r)\,\left(\psi_{\ell_{1},n}^{2}(r)-1\right)\,dr\right)_{n\geq 1},\quad\left(\int_{0}^{1}\zeta(r)\,\left(\psi_{\ell_{2},n}^{2}(r)-1\right)\,dr\right)_{n\geq 1}\right),$
where the eigenfunctions $\psi_{\ell,n}(r)$ are given, up to a normalization [^2] As mentioned earlier, we shall omit this inessential normalization., by
$\psi_{\ell,n}(r):=\sqrt{\frac{\pi r}{2}}\,J_{\nu}(j_{\nu,n}r),\quad\nu=\ell+\tfrac{1}{2},\quad n\geq 1.$
We now state the linearized uniqueness result in terms of the spectral map and its differential.
Theorem 2.2 (Injectivity of the differential of the spectral map).
Let $\ell_{1}\neq\ell_{2}$ be two distinct non-negative integers, and consider the spectral map
$\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\to\mathbb{R}\times\ell_{\mathbb{R}}^{2}(\mathbb{N})\times\ell_{\mathbb{R}}^{2}(\mathbb{N}).$
Then the Fréchet differential at the zero potential $q=0$
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}:L^{2}(0,1)\to\mathbb{R}\times\ell_{\mathbb{R}}^{2}(\mathbb{N})\times\ell_{\mathbb{R}}^{2}(\mathbb{N}),$
is injective, provided that $(\ell_{1},\ell_{2})=(0,1)$ or $(\ell_{1},\ell_{2})=(0,2)$.
The proof of Theorem 2.2 is given in Sections 7–8.
As a byproduct of our analysis we obtain the following theorem in the case $\ell_{1}=0$ and arbitrary $\ell_{2}$.
Theorem 2.3 (Finite-dimensional kernel).
Let $\ell_{2}\geq 1$. Consider
$\mathcal{S}_{0,\ell_{2}}:L^{2}(0,1)\longrightarrow\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N}).$
Then the Fréchet differential of $\mathcal{S}_{0,\ell_{2}}$ at the zero potential $q=0$ has a finite-dimensional kernel.
Sketch of proof.
Our approach shows that any element of this kernel satisfies a homogeneous linear ODE on $(0,1)$, whose order and coefficients depend only on $\ell_{2}$. The cases $(\ell_{1},\ell_{2})=(0,1)$ and $(0,2)$ are treated in detail in this paper, and the extension to arbitrary $\ell_{2}$ follows without additional difficulty. Consequently, the kernel is spanned by finitely many fundamental solutions and is therefore finite-dimensional. ∎
This recovers the case $(\ell_{1},\ell_{2})=(0,1)$ proved in [25], and is consistent with [4] when $\ell_{1}-\ell_{2}$ is odd. We conjecture that the same conclusion holds for all pairs of distinct nonnegative integers $\ell_{1}\neq\ell_{2}$.
2.2 Proof of Theorem 1.3
In this subsection, we show how Theorem 2.2 implies Theorem 1.3. First, let us decompose $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ in two parts
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}=A_{\ell_{1},\ell_{2}}+K_{\ell_{1},\ell_{2}}$
where, denoting for any integer $\ell$, $\Phi_{\ell}(x):=\frac{\pi}{2}xJ_{\ell+\frac{1}{2}}(x)^{2}$,
$A_{\ell_{1},\ell_{2}}(\zeta)=\Bigg(\int_{0}^{1}\zeta(r)\,dr,$ $\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi r\right)-1\Bigr)\,dr\right)_{n\geq 1},$
$\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi r\right)-1\Bigr)\,dr\right)_{n\geq 1}\Bigg).$ (2.2)
Then, we have:
$K_{\ell_{1},\ell_{2}}(\zeta)=\Bigg($  $\,0,\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{1}}(j_{\ell_{1}+\frac{1}{2},n}\pi r)-\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi r\right)\Bigr)\,dr\right)_{n\geq 1},$
$\left(\int_{0}^{1}\zeta(r)\,\Bigl(\Phi_{\ell_{2}}(j_{\ell_{2}+\frac{1}{2},n}\pi r)-\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi r\right)\Bigr)\,dr\right)_{n\geq 1}\Bigg).$ (2.3)
Since $|\Phi^{\prime}_{\ell}(x)|\leq C_{\ell}$ uniformly on $\mathbb{R}$ and $\sum_{n}|j_{\ell+\frac{1}{2},n}-(n+\frac{\ell}{2})|^{2}<\infty$ for any integer value of $\ell$, we get
$\sum_{n}\|\Phi_{\ell}(j_{\ell,n}\pi\cdot)-\Phi_{\ell}((n+\frac{\ell}{2})\pi\cdot)\|_{L^{2}(0,1)}^{2}<\infty$
and thus, by Cauchy–Schwarz inequality, $K_{\ell_{1},\ell_{2}}$ is a Hilbert–Schmidt (hence compact) operator from $L^{2}(0,1)$ into $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$. Now from ([23], Corollary 5.2) and Theorem A.2 in the appendix of this paper), we learn that, for $(\ell_{1},\ell_{2})=(0,1)$ or $(\ell_{1},\ell_{2})=(0,2)$,
$\left(1,\ \left(\Phi_{\ell_{1}}((n+\frac{\ell_{1}}{2})\pi\cdot)\right),\ \left(\Phi_{\ell_{2}}((n+\frac{\ell_{2}}{2})\pi\cdot)\right)\right)$
is a basis of $L^{2}(0,1)$ which in turn implies that $A_{\ell_{1},\ell_{2}}$ is an isomorphism from $L^{2}(0,1)$ onto $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$. Therefore by Theorem 2.2 and the Fredholm alternative, $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ is also an isomorphism from $L^{2}(0,1)$ onto $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$. Now, applying the inverse function theorem, we obtain Theorem 1.3.
3 The Kneser–Sommerfeld formula
The so-called Kneser–Sommerfeld expansion is a classical identity involving Bessel functions and their zeros. It appears in Watson’s treatise [30] (Section 15.42), where it is stated as a Fourier–Bessel series identity. However, as first noted by Buchholz in 1947 [2], and more recently clarified by Martin [15], the formula given in Watson is incorrect: it is missing an integral term on the right-hand side.
This correction was independently rediscovered in 1981 by Kobayashi [10], who analyzed the same identity and showed that Watson’s version could not be valid in general. However, a proper historical credit for the correction belongs to Buchholz, whose 1947 paper already pointed out the failure and provided a corrected structure.
In his detailed analysis, Martin [15] confirms that a different version - found in earlier works by Kneser [11] for $\nu=0$, Sommerfeld [26] for integer values of $\nu$, and Carslaw [3] - is in fact the correct identity. Surprisingly, Watson does not include this valid form, despite its mathematical consistency and symmetry.
The corrected version of the Kneser–Sommerfeld expansion reads:
$\sum_{n=1}^{\infty}\frac{J_{\nu}(xj_{\nu,n})J_{\nu}(Xj_{\nu,n})}{(z^{2}-j_{\nu,n}^{2})\left[J^{\prime}_{\nu}(j_{\nu,n})\right]^{2}}=\frac{\pi}{4J_{\nu}(z)}J_{\nu}(xz)\left[J_{\nu}(z)Y_{\nu}(Xz)-Y_{\nu}(z)J_{\nu}(Xz)\right],$ (3.1)
valid for $0\leq x\leq X\leq 1$, $\nu\in\mathbb{R}$, $z\not=j_{\nu,n}$, where $J_{\nu}$ and $Y_{\nu}$ are the Bessel functions of the first and second kinds, respectively, and $j_{\nu,n}$ denotes the $n$-th positive zero of $J_{\nu}$.
The identity (3.1) plays a fundamental role in the analysis of our inverse spectral problem.
3.1 An integral Green identity
In the next subsection, we will present a first completeness theorem for the inverse problem under consideration. The argument relies solely on the moment formula (3.4) together with the Müntz–Szász theorem, and yields uniqueness under a rather restrictive vanishing condition on the moments.
So, we begin by considering a function $\zeta\in L^{2}(0,1)$ that is orthogonal to the family of squared eigenfunctions associated with a fixed angular momentum $\ell\in\mathbb{N}$ and vanishing potential. That is,
$\int_{0}^{1}x\,\zeta(x)\left[J_{\nu}(j_{\nu,n}x)\right]^{2}dx=0,\qquad\text{for all }n\geq 1,$ (3.2)
where $\nu=\ell+\tfrac{1}{2}$ and $j_{\nu,n}$ denotes the $n$-th positive zero of the Bessel function $J_{\nu}$.
Our goal is to exploit the structure of the family $\left\{\left[J_{\nu}(j_{\nu,n}x)\right]^{2}\right\}_{n\geq 1}$ using the Kneser–Sommerfeld expansion (3.1), and to derive an additional integral identity satisfied by $\zeta$.
To that end, we consider a weighted sum of the orthogonality relations (3.2) over all $n\geq 1$. By exchanging the order of summation and integration (justified, for instance, by Fubini’s theorem), we obtain for all $z$ not equal to any zero $j_{\nu,n}$:
$\int_{0}^{1}x\,\zeta(x)\,J_{\nu}(xz)\left[J_{\nu}(z)\,Y_{\nu}(xz)-Y_{\nu}(z)\,J_{\nu}(xz)\right]dx=0.$ (3.3)
The expression inside the integral corresponds, up to a multiplicative constant, to the kernel appearing in the corrected Kneser–Sommerfeld expansion on the diagonal $x=X$ multiplied by $J_{\nu}(z)$. By a standard continuity argument, this identity extends to all complex values of $z$.
3.2 A first completeness result
In this subsection we prove Theorem 1.2, the completeness result stated in the introduction. Our strategy is to establish first an integral identity, which will then allow us to deduce a moment condition. Combining this with the classical Müntz–Szász theorem ultimately leads to the desired completeness result, which appears as a direct consequence of the final theorem in this section.
Recall that, as $x\to 0$, the Bessel functions admit the following asymptotic behavior (see [14]):
$J_{\nu}(x)\sim\frac{1}{\Gamma(\nu+1)}\left(\frac{x}{2}\right)^{\nu},\quad Y_{\nu}(x)\sim-\frac{\Gamma(\nu)}{\pi}\left(\frac{2}{x}\right)^{\nu}\quad\text{for }\nu>0.$
Thus, taking the limit $z\to 0$ in (3.3), we obtain the following closed-form expression:
$\int_{0}^{1}x\ \zeta(x)\,\left(1-x^{2\nu}\right)dx=0.$ (3.4)
We can now state a first completeness result based on the classical Müntz–Szász theorem ([16, 17, 27]). It is worth emphasizing that, in the following theorem, no zero-mean condition is required for the function $\zeta$.
Theorem 3.1 (A completeness result).
Let $\zeta\in L^{2}(0,1)$ be a real-valued function satisfying (3.2) for an infinite increasing sequence $\nu_{k}=\{\ell_{k}+\tfrac{1}{2}\}$ of positive half-integers such that
$\sum_{k=1}^{\infty}\frac{1}{\ell_{k}}=+\infty.$
Then $\zeta=0$ almost everywhere in $(0,1)$.
Proof.
Let $\{\ell_{k}\}\subset\mathbb{N}^{*}$ be the sequence of positive integers appearing above, with $\sum_{k=1}^{\infty}\frac{1}{\ell_{k}}=+\infty$, and apply identity (3.4) with $\nu_{k}=\ell_{k}+\tfrac{1}{2}$. This yields, as $k\to+\infty$,
$\int_{0}^{1}x\,\zeta(x)\,dx=0.$
Plugging this back into (3.4), we obtain the moment identities
$\int_{0}^{1}\zeta(x)\,x^{2\ell_{k}+2}\,dx=0\quad\text{for all }k\in\mathbb{N}.$
By the classical Müntz–Szász theorem, this implies that $\zeta=0$ almost everywhere in $(0,1)$. ∎
As explained at the beginning of this section, Theorem 3.1 immediately yields Theorem 1.2, thereby completing the proof of the first completeness result.
4 Transformation Operators
Transformation operators play a central role in our inverse spectral analysis. For the case $\ell=1$, such operators were first introduced by Guillot and Ralston [9] in their study of radial Schrödinger operators. These operators were also used by C. Shubin Christ in [25]. Carlson and Shubin [4] extended this construction to more general settings, and Rundell and Sacks [23] developed a systematic framework valid for all integer orders $\ell\geq 1$. Their method rewrites inner products involving Bessel kernels in terms of trigonometric ones by means of suitable index-reduction operators.
We recall below two key lemmas due to Rundell and Sacks [23] (also used by Serier [24]). These lemmas provide the basic tools for reducing the Bessel case to the trigonometric one and will be crucial for the fine analysis of isospectral sets developed later.
Notation 4.1.
Let $\nu=\ell+\tfrac{1}{2}$ for $\ell\geq 0$ and $x\in[0,1]$. Define
$\Phi_{\ell}(x)\;=\;\frac{\pi x}{2}\,J_{\nu}(x)^{2},\qquad\Psi_{\ell}(x)\;=\;-\,\frac{\pi x}{2}\,J_{\nu}(x)\,Y_{\nu}(x),$
where $J_{\nu}$, $Y_{\nu}$ are the ordinary Bessel functions.
Lemma 4.1.
[Rundell–Sacks [23]] For each integer $\ell\geq 1$ and $x\in(0,1)$, define the index-reduction operator
$S_{\ell}[f](x)\;=\;f(x)-4\ell\,x^{2\ell-1}\int_{x}^{1}t^{-2\ell}\,f(t)\,dt.$ (4.1)
Then:
  1. (i)
    $S_{\ell}$ is bounded on $L^{2}(0,1)$.
  2. (ii)
    Its Hilbert adjoint is
    $S_{\ell}^{*}[g](x)\;=\;g(x)-\frac{4\ell}{x^{2\ell}}\int_{0}^{x}t^{2\ell-1}\,g(t)\,dt.$ (4.2)
  3. (iii)
    It induces a Banach isomorphism
    $L^{2}(0,1)\;\xrightarrow{\sim}\;\bigl\{x\mapsto x^{2\ell}\bigr\}^{\perp},$
    with inverse
    $A_{\ell}[g](x)\;=\;g(x)-\frac{4\ell}{x^{2\ell+1}}\int_{0}^{x}t^{2\ell}\,g(t)\,dt.$ (4.3)
  4. (iv)
    If $g=S_{\ell}[f]$, then $f$ and $g$ satisfy
    $f^{(2\ell)}(x)+\frac{4\ell}{x}\,f^{(2\ell-1)}(x)\;=\;g^{(2\ell)}(x).$ (4.4)
  5. (v)
    The operators commute pairwise:
    $S_{\ell}\,S_{m}\;=\;S_{m}\,S_{\ell},\quad\forall\,\ell,m\in\mathbb{N}.$
  6. (vi)
    The functions $\Phi_{\ell}$ and $\Psi_{\ell}$ satisfy
    $\Phi_{\ell}$ $=-\,S_{\ell}^{*}\bigl[\Phi_{\ell-1}\bigr],$ $\Psi_{\ell}$ $=-\,S_{\ell}^{*}\bigl[\Psi_{\ell-1}\bigr],$ (4.5)
    $\Phi_{\ell}^{\prime}$ $=-\,A_{\ell}\bigl[\Phi_{\ell-1}^{\prime}\bigr],$ $\Psi_{\ell}^{\prime}$ $=-\,A_{\ell}\bigl[\Psi_{\ell-1}^{\prime}\bigr].$ (4.6)
We will also need the following complementary result (see Lemma 3.4 in [23]).
Lemma 4.2.
If for $\ell\geq 1$, $f$ and $g\in L^{2}(0,1)$, it holds $g=S_{\ell}^{*}[f]$ then, in $\mathcal{D}^{\prime}(0,1)$, we have
$g^{(2\ell+1)}(x)+\frac{4\ell}{x}g^{(2\ell)}(x)=f^{(2\ell+1)}(x).$
Proof.
We follow the strategy of [23] for the proof of (4.4). We start from
$g(x)=f(x)-4\ell x^{-2\ell}\int_{0}^{x}s^{2\ell-1}f(s)\,ds$ (4.7)
and differentiate once for getting
$g^{\prime}(x)=f^{\prime}(x)+8\ell^{2}x^{-2\ell-1}\int_{0}^{x}s^{2\ell-1}f(s)\,ds-\frac{4\ell}{x}f(x).$ (4.8)
[…]
$\frac{2\ell}{x}g(x)+g^{\prime}(x)=\frac{2\ell}{x}f(x)+f^{\prime}(x)-\frac{4\ell}{x}f(x).$
This leads to
$2\ell g(x)+xg^{\prime}(x)=-2\ell f(x)+xf^{\prime}(x).$
Differentiating again, we get
$(2\ell+1)g^{\prime}(x)+xg^{\prime\prime}(x)=(-2\ell+1)f^{\prime}(x)+xf^{\prime\prime}(x)$
and iterating $k$ times we get
$(2\ell+k)g^{(k)}(x)+xg^{(k+1)}(x)=(-2\ell+k)f^{(k)}(x)+xf^{(k+1)}(x).$
Taking $k=2\ell$ and dividing by $x$ achieves the proof of the lemma. ∎
We now consider the composite operator $T_{\ell}$, obtained by composing the index-reduction operators $S_{1},\dots,S_{\ell}$, which carries Bessel kernels to trigonometric ones.
Lemma 4.3.
[Rundell–Sacks [23], Serier [24]] Let us define $T_{\ell}$ by $T_{1}=S_{1}$ and for $\ell\geq 2$ by
$T_{\ell}\;=\;(-1)^{\ell+1}\,S_{\ell}\,S_{\ell-1}\cdots S_{1}.$
Then:
  1. (i)
    $T_{\ell}$ is a bounded and injective operator on $L^{2}(0,1)$. For any $\zeta\in L^{2}(0,1)$ and $z\in\mathbb{C}$,
    $\int_{0}^{1}\bigl[2\,\Phi_{\ell}(zt)-1\bigr]\,\zeta(t)\,dt$ $=\int_{0}^{1}\cos(2zt)\,T_{\ell}[\zeta](t)\,dt,$ (4.9)
    $\int_{0}^{1}\Psi_{\ell}(zt)\,\zeta(t)\,dt$ $=-\tfrac{1}{2}\int_{0}^{1}\sin(2zt)\,T_{\ell}[\zeta](t)\,dt.$ (4.10)
  2. (ii)
    Its adjoint satisfies
    $T_{\ell}^{*}[\cos(2zx)]\;=\;2\,\Phi_{\ell}(zx)-1,\qquad T_{\ell}^{*}\bigl[-\tfrac{1}{2}\sin(2zx)\bigr]\;=\;\Psi_{\ell}(zx),$
    with $\ker(T_{\ell}^{*})=\mathrm{span}\{x^{2},x^{4},\dots,x^{2\ell}\}.$
  3. (iii)
    $T_{\ell}$ provides a Banach isomorphism
    $L^{2}(0,1)\;\xrightarrow{\sim}\;\bigl(\ker T_{\ell}^{*}\bigr)^{\perp},$
    whose inverse is
    $B_{\ell}[f]\;=\;(-1)^{\ell+1}\,A_{\ell}\,A_{\ell-1}\cdots A_{1}[f].$
    Moreover,
    $\Phi_{\ell}^{\prime}(zx)\;=\;B_{\ell}[-\sin(2zx)],\qquad\Psi_{\ell}^{\prime}(zx)\;=\;B_{\ell}[-\cos(2zx)].$
5 Green’s identity via transformation operators
In this section, we shall employ the transformation operators $T_{\ell}$ to reformulate relation (3.3) in a more convenient form. First, it is straightforward to verify that,
$J_{\nu}(xz)^{2}=\frac{2}{\pi\,z\,x}\,\Phi_{\ell}(zx),\qquad J_{\nu}(xz)\,Y_{\nu}(xz)=-\frac{2}{\pi\,z\,x}\,\Psi_{\ell}(zx).$
Substituting these into (3.3) gives
$J_{\nu}(z)\,\int_{0}^{1}\Psi_{\ell}(zx)\,\zeta(x)\,dx\;+\;Y_{\nu}(z)\,\int_{0}^{1}\Phi_{\ell}(zx)\,\zeta(x)\,dx\;=\;0.$ (5.1)
Moreover, for functions $\zeta$ orthogonal to the constants on $[0,1]$, i.e. satisfying $\int_{0}^{1}\zeta(x)\,dx=0,$ it follows from (5.1) that
$2\,J_{\nu}(z)\,\int_{0}^{1}\Psi_{\ell}(zx)\,\zeta(x)\,dx\;+\;Y_{\nu}(z)\,\int_{0}^{1}\bigl(2\,\Phi_{\ell}(zx)-1\bigr)\,\zeta(x)\,dx\;=\;0\,.$ (5.2)
We now invoke Lemma 4.3(i) together with the identity
$Y_{\nu}(x)=(-1)^{\,\ell+1}J_{-\nu}(x)$
to deduce that, for such $\zeta$, (5.2) can be equivalently written as
$J_{\nu}(z)\int_{0}^{1}T_{\ell}[\zeta](x)\,\sin(2zx)\,dx\;+\;(-1)^{\,\ell}\,J_{-\nu}(z)\int_{0}^{1}T_{\ell}[\zeta](x)\,\cos(2zx)\,dx\;=\;0\,.$ (5.3)
We will now employ the half-integer Bessel functions and their associated polynomials $P_{\ell}$ and $Q_{\ell}$ (see [1, 10.1.19–20]). For $\ell=0,1,2,\dots$ and $z\in\mathbb{C}$ ,
$J_{\ell+\frac{1}{2}}(z)$ $=\sqrt{\frac{2}{\pi z}}\,\Bigl[P_{\ell}\!\bigl(\tfrac{1}{z}\bigr)\,\sin z\;-\;Q_{\ell-1}\!\bigl(\tfrac{1}{z}\bigr)\,\cos z\Bigr],$ (5.4)
$J_{-\,\ell-\frac{1}{2}}(z)$ $=(-1)^{\,\ell}\sqrt{\frac{2}{\pi z}}\,\Bigl[P_{\ell}\!\bigl(\tfrac{1}{z}\bigr)\,\cos z\;+\;Q_{\ell-1}\!\bigl(\tfrac{1}{z}\bigr)\,\sin z\Bigr]\,.$ (5.5)
The polynomials $P_{\ell}(t)$ and $Q_{\ell}(t)$, each of degree $\ell$ in $t$, satisfy the three-term recursion
$P_{\ell+1}(t)$ $=(2\ell+1)\,t\,P_{\ell}(t)\;-\;P_{\ell-1}(t)\,,\qquad\ell\geq 1,$ (5.6)
$Q_{\ell+1}(t)$ $=(2\ell+3)\,t\,Q_{\ell}(t)\;-\;Q_{\ell-1}(t)\,,\qquad\ell\geq 0,$ (5.7)
with initial conditions
$P_{0}(t)=1\,,\quad P_{1}(t)=t\,,\qquad Q_{-1}(t)=0\,,\quad Q_{0}(t)=1\,.$
Note that $P_{\ell}(t)$ and $Q_{\ell}(t)$ are even polynomials if $\ell$ is even, and odd polynomials if $\ell$ is odd.
As a concrete example, the simplest half-integer Bessel functions are
$J_{\tfrac{1}{2}}(z)=\sqrt{\frac{2}{\pi z}}\;\sin z\,,\qquad J_{-\tfrac{1}{2}}(z)=\sqrt{\frac{2}{\pi z}}\;\cos z\,.$ (5.8)
The next case corresponds to order $\pm\tfrac{3}{2}$:
$J_{\tfrac{3}{2}}(z)=\sqrt{\frac{2}{\pi z}}\left(\frac{\sin z}{z}-\cos z\right),\qquad J_{-\tfrac{3}{2}}(z)=\sqrt{\frac{2}{\pi z}}\left(-\frac{\cos z}{z}-\sin z\right).$ (5.9)
The first few polynomials $P_{\ell}(t)$ and $Q_{\ell}(t)$ generated by the recursion relations are:
$P_{0}(t)$ $=1\,$, $Q_{-1}(t)$ $=0\,$, (5.10)
$P_{1}(t)$ $=t\,$, $Q_{0}(t)$ $=1\,$,
$P_{2}(t)$ $=3t^{2}-1\,$, $Q_{1}(t)$ $=3t\,$,
[…]
From the foregoing we obtain the following:
Proposition 5.1.
Let $\zeta\in L^{2}(0,1)$ be orthogonal to the constants and suppose further that for $\nu=\ell+\frac{1}{2}$,
$\int_{0}^{1}x\,\zeta(x)\,J_{\nu}\bigl(j_{\nu,n}\,x\bigr)^{2}\,dx\;=\;0\,,\qquad\text{for all }n\geq 1\,.$
Then for every $z\in\mathbb{C}$ and every integer $\ell\geq 1$ ,
[…]
Proof.
The result follows immediately from (5.3), (5.4) and (5.5), by a straightforward application of elementary trigonometric identities. ∎
Remark 5.2.
In the foregoing we have relied on the transformation operators $S_{\ell}$ and $T_{\ell}$ defined for $\ell\geq 1$. For the case $\ell=0$, (i.e $\nu=\frac{1}{2}$), one obtains an analogous reduction by using the explicit formulas (5.8). It follows that
$\Phi_{0}(x)=\frac{1-\cos(2x)}{2},\qquad\Psi_{0}(x)=\tfrac{1}{2}\sin(2x).$
Hence, under the hypotheses of Proposition 5.1, it follows from (5.2) that
$\int_{0}^{1}\zeta(x)\,\cos\bigl(z(2x-1)\bigr)\,dx=0\quad\text{for all }z\in\mathbb{C}.$ (5.12)
In other words, if one conventionally sets $T_{0}=\mathrm{Id}$, then relation (5.11) also extends to the case $\ell=0$.
Remark 5.3.
Note that if one multiplies (5.11) by $z^{\ell}$ and then sets $z=0$, it follows that for $\ell\geq 1$,
$\int_{0}^{1}T_{\ell}[\zeta](x)\,dx=0.$
6 From Green’s integral identities to differential relations
Let $\zeta\in L^{2}(0,1)$ be orthogonal to the constants and satisfy the hypotheses of Proposition 5.1. Then, for each integer $\ell\geq 0$, we shall show that the integral identity (5.11) gives rise to a differential relation satisfied by the transformed function $T_{\ell}[\zeta]$.
We introduce the reflection symmetry about the midpoint $x=\tfrac{1}{2}$, defined by the involutive map
$\sigma(x):=1-x.$
This transformation naturally arises in several parts of the analysis, particularly when exploiting parity properties. Throughout the paper, we refer to this symmetry simply as $\sigma$.
6.1 The case $\ell=0$
By standard Fourier series theory, the relation (5.12) implies that $\zeta$ is an odd function with respect to the midpoint $x_{0}=\tfrac{1}{2}$, i.e.
$\zeta(x)=-\,\zeta\bigl(\sigma(x)\bigr)\quad\text{for almost every }x\in[0,1].$
6.2 The case $\ell=1$
According to (5.10), and setting $g=T_{1}(\zeta)=S_{1}(\zeta)$, equation (5.11) for $\ell=1$ can be written as
$\int_{0}^{1}g(x)\,\Bigl[\tfrac{1}{z}\cos\bigl(z(2x-1)\bigr)-\sin\bigl(z(2x-1)\bigr)\Bigr]\,dx=0.$ (6.1)
Now, we define
$G(x)=\int_{0}^{x}g(t)\,\mathrm{d}t.$
By Remark 5.3, $G(1)=0$, so using integration by parts, we deduce
$\int_{0}^{1}\bigl[\,2G(x)-g(x)\bigr]\sin\bigl(z(2x-1)\bigr)\,\mathrm{d}x=0.$ (6.2)
In other words, 2G(x)g(x)2G(x)-g(x) is an even function with respect to $x=\tfrac{1}{2}$.
6.3 The case $\ell\geq 2$
Before stating the result, we briefly recall the notion of symmetry for distributions.
Definition 6.1.
Let $\sigma(x):=1-x$ denote the reflection with respect to the midpoint of the interval $[0,1]$. This map induces a natural action on distributions $T\in\mathcal{D}^{\prime}(0,1)$ via pullback, defined by duality:
$\langle\sigma_{*}T,\varphi\rangle:=\langle T,\varphi\circ\sigma\rangle,\qquad\text{for all }\varphi\in\mathcal{D}(0,1).$
We say that a distribution $T\in\mathcal{D}^{\prime}(0,1)$ is even if $\sigma_{*}T=T$, and odd if $\sigma_{*}T=-T$.
Remark 6.2.
There is a classical characterization of even and odd distributions on the interval $(0,1)$ using Fourier series adapted to the symmetry with respect to the midpoint $x=\tfrac{1}{2}$.
Consider the orthonormal basis of $L^{2}(0,1)$ formed by the functions
$\phi_{n}(x):=\sqrt{2}\cos\big(\pi n(2x-1)\big),\quad n\geq 0\mbox{ and }\psi_{n}(x):=\sqrt{2}\sin\big(\pi n(2x-1)\big),\quad n\geq 1.$
Each $\phi_{n}$ is even with respect to $x=\tfrac{1}{2}$, while each $\psi_{n}$ is odd.
Then, a distribution $T\in\mathcal{D}^{\prime}(0,1)$ is odd with respect to $x=\tfrac{1}{2}$ if and only if
$\langle T,\phi_{n}\rangle=0\qquad\text{for all }n\geq 1.$
Similarly, $T$ is even if and only if $\langle T,\psi_{n}\rangle=0$ for all $n\geq 1$.
Equivalently, a distribution $T\in\mathcal{D}^{\prime}(0,1)$ is odd with respect to the midpoint $x=\tfrac{1}{2}$ if
$\langle T,\cos\big(z(2x-1)\big)\rangle=0\qquad\text{for all }z\in\mathbb{C},$
and it is even if
$\langle T,\sin\big(z(2x-1)\big)\rangle=0\qquad\text{for all }z\in\mathbb{C}.$
We now introduce the following transformed polynomials:
$\widetilde{P}_{\ell}(z):=z^{\ell}P_{\ell}\left(\frac{1}{z}\right),\qquad\widetilde{Q}_{\ell-1}(z):=z^{\ell}Q_{\ell-1}\left(\frac{1}{z}\right),$
where $P_{\ell}(t)$ and $Q_{\ell-1}(t)$ are as previously defined.
By construction, $\widetilde{P}_{\ell}(z)$ is an even polynomial for all $\ell$, while $\widetilde{Q}_{\ell-1}(z)$ is always odd, regardless of the parity of $\ell$. We list below the first few examples:
$\widetilde{P}_{1}(z)$ $=1\,$, $\qquad\widetilde{Q}_{0}(z)$ $=z\,$,
$\widetilde{P}_{2}(z)$ $=3-z^{2}\,$, $\qquad\widetilde{Q}_{1}(z)$ $=3z\,$,
$\widetilde{P}_{3}(z)$ $=15-6z^{2}\,$, $\qquad\widetilde{Q}_{2}(z)$ $=15z-z^{3}\,.$
We also introduce the differential operator
$\widetilde{D}:=\frac{1}{2i}\frac{d}{dx},$
so that for any polynomial $R(z)=\sum_{k}a_{k}z^{k}$, the corresponding operator R(D~)R($\widetilde${D}) acts on a function $f$ via
$R(\widetilde{D})f(x)=\sum_{k}a_{k}(\widetilde{D}^{k}f)(x)\,.$
We now proceed by multiplying the identity (5.11) by $z^{\ell}$. Under the assumptions of Lemma 5.1, we obtain the following identity.
Lemma 6.3.
Let $\zeta\in L^{2}(0,1)$ satisfy the hypotheses of Lemma 5.1, and let $g:=T_{\ell}[\zeta]$. Then, one has for all $z\in\mathbb{C}$,
$\int_{0}^{1}g(x)\,\cdot\,\Bigl[\widetilde{P}_{\ell}(\widetilde{D})+i\,\widetilde{Q}_{\ell-1}(\widetilde{D})\Bigr]\cos(z(2x-1))\,dx=0\,.$
Remark 6.4.
As a consistency check, let us consider the case $\ell=1$. In this case, one has
$\widetilde{P}_{1}(\widetilde{D})+i\,\widetilde{Q}_{0}(\widetilde{D})=1+\frac{1}{2}\frac{d}{dx}\,.$
Thus, we obtain:
$\int_{0}^{1}g(x)\left(1+\frac{1}{2}\frac{d}{dx}\right)\cos\big(z(2x-1)\big)\,dx=\int_{0}^{1}g(x)\,\left[\cos\big(z(2x-1)\big)-z\sin\big(z(2x-1)\big)\right]\,dx=0\,$,
as already observed in (6.1).
In particular, Lemma 6.3 implies that the distribution
$\left(\widetilde{P}_{\ell}(\widetilde{D})-i\,\widetilde{Q}_{\ell-1}(\widetilde{D})\right)g(x)$
is odd with respect to the midpoint $x=\tfrac{1}{2}$. This follows from the fact that $\widetilde{P}_{\ell}$ is an even polynomial, while $\widetilde{Q}_{\ell-1}$ is odd.
We are now left with understanding more precisely the differential operator
[…]
Using the recursion relations satisfied by the polynomials $P_{\ell}$ and $Q_{\ell}$, and introducing
$\widetilde{A}_{\ell}(z):=\widetilde{P}_{\ell}(z)-i\,\widetilde{Q}_{\ell-1}(z),$
we obtain the following lemma.
Lemma 6.5.
For every integer $\ell\geq 1$, the polynomials $\widetilde{A}_{\ell}(z)$ satisfy the recursion relation
$\widetilde{A}_{\ell+1}(z)-(2\ell+1)\widetilde{A}_{\ell}(z)+z^{2}\widetilde{A}_{\ell-1}(z)=0\,$,
with initial values
$\widetilde{A}_{0}(z)=1\,,\qquad\widetilde{A}_{1}(z)=1-iz\,.$
At this stage, the polynomials $\widetilde{A}_{\ell}(z)$ have complex coefficients. To simplify the situation, we set $A_{\ell}(t)=\tilde{A}_{\ell}(\frac{t}{2i})$. We obtain the following result, which also holds for $\ell=0$ and $\ell=1$:
Theorem 6.6.
Let $\{A_{\ell}(t)\}_{\ell\in\mathbb{N}}$ be the sequence of polynomials defined recursively by
$A_{0}(t)=1\,,\qquad A_{1}(t)=1-\frac{t}{2}\,$,
and for all $\ell\geq 1$ ,
$A_{\ell+1}(t)=(2\ell+1)\,A_{\ell}(t)+\frac{t^{2}}{4}\,A_{\ell-1}(t)\,.$
Suppose that $\zeta$ satisfies the hypotheses of Lemma 5.1. Then, in the sense of distributions, the function
$A_{\ell}\bigl(D\bigr)\bigl[T_{\ell}[\zeta]\bigr]\,$,
where $D=\frac{d}{dx}$, is odd with respect to the midpoint $x=\tfrac{1}{2}$ .
Remark 6.7.
The next polynomials in the sequence are given by
$A_{2}(t)\;=\;\tfrac{1}{4}t^{2}-\tfrac{3}{2}t+3,\qquad A_{3}(t)\;=\;-\tfrac{1}{8}t^{3}+\tfrac{3}{2}t^{2}-\tfrac{15}{2}t+15.$
7 Uniqueness result in the case $\ell=0$ and $\ell=1$
In this section, we assume that $\zeta$ satisfies the hypotheses of Propostion 5.1 for $\ell=0$ and $\ell=1$ . For $\ell=0$ , this simply implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$. For $\ell=1$ , setting $g:=T_{1}[\zeta]$, the result means that the function $g^{\prime}(x)+2g(x)$ is odd in the sense of distributions.
The goal is to reduce the problem to the study of an ordinary differential equation involving $\zeta$ only. According to Lemma 4.1 (vi), applied with $\ell=1$, and since $g=S_{1}[\zeta]$, we have:
$g^{\prime\prime}(x)=\zeta^{\prime\prime}(x)+\frac{4}{x}\,\zeta^{\prime}(x)\,.$ (7.1)
Differentiating both sides gives:
$g^{(3)}(x)=\zeta^{(3)}(x)+\frac{4}{x}\zeta^{\prime\prime}(x)-\frac{4}{x^{2}}\zeta^{\prime}(x)\,.$ (7.2)
Since $g^{\prime}(x)+2g(x)$ is odd, it follows that $g^{(3)}(x)+2g^{\prime\prime}(x)$ is also odd in the sense of distributions. Using equations (7.1) and (7.2), we deduce that the following expression is odd:
$f(x):=-\zeta^{(3)}(x)+\left(2-\frac{4}{x}\right)\zeta^{\prime\prime}(x)+\left(\frac{4}{x^{2}}+\frac{8}{x}\right)\zeta^{\prime}(x)\,.$
Letting $y:=\zeta^{\prime}(x)$, we rewrite this as
$f(x)=-y^{\prime\prime}(x)+\left(2-\frac{4}{x}\right)y^{\prime}(x)+\left(\frac{4}{x^{2}}+\frac{8}{x}\right)y(x)\,$, (7.3)
which is odd with respect to $x=\tfrac{1}{2}$, that is satisfies, f(x)=f(1x)f(x)=-f(1-x) .
Taking into account the parity of $y$, and after elementary calculations, we arrive at the following lemma:
Lemma 7.1.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Lemma 5.1 for $\ell=0$ and $\ell=1$, and set $y(x):=\zeta^{\prime}(x)$. Then $y$ satisfies the differential equation
$y^{\prime\prime}(x)+\left(\frac{2}{x}-\frac{2}{1-x}\right)y^{\prime}(x)-\left(\frac{2}{x^{2}}+\frac{2}{(1-x)^{2}}+\frac{4}{x}+\frac{4}{1-x}\right)y(x)=0$ (7.4)
in the sense of distributions on $(0,1)$.
Now, we seek symmetric solutions of the differential equation (7.4). The equation is weakly singular at both endpoints $x=0$ and $x=1$ ([28], p. 236, or [29]), and we begin by analyzing the behavior near $x=0$. The indicial equation associated with the singularity at zero is
$\rho(\rho-1)+2\rho-2=0,$
which has roots $\rho=1$ and $\rho=-2$. There is a basis of solutions near zero with one asymptotic to $x^{-2}$ and the other asymptotic to $x$. The solution which is asymptotic to $x^{-2}$ is to exclude since this would lead to a $\zeta$ which is not in $L^{2}(0,1)$. Consequently, our solution is proportional to the second one and satisfies $y(0)=0$.
A similar analysis near $x=1$ leads to the same conclusion: the admissible solution must satisfy $y(1)=0$. Hence, we restrict ourselves to functions $y\in H_{0}^{1}(0,1)$ satisfying homogeneous Dirichlet boundary conditions.
Since $f(x)$ given in (7.3) is odd and $y(x)$ is even, we have
[…]
A straightforward integration by parts, using the boundary conditions $y(0)=y(1)=0$, yields
$0=\int_{0}^{1}f(x)\,y(x)\,dx\;=\;\int_{0}^{1}\bigl(y^{\prime}(x)\bigr)^{2}\,dx\;+\;\int_{0}^{1}\left(\frac{2}{x^{2}}+\frac{8}{x}\right)\,y^{2}(x)\,dx.$ (7.5)
It follows that $y(x)=0$ almost everywhere. We conclude that $\zeta$ must be constant. But since $\zeta$ is odd with respect to $x=\tfrac{1}{2}$, we get $\zeta=0$.
We conclude this section with the following completeness result:
Theorem 7.2.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Proposition 5.1 for $\ell=0$ and $\ell=1$. Then $\zeta=0$ almost everywhere on $(0,1)$.
In particular, this implies that the spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,1)$ is injective.
8 Uniqueness result in the case $\ell=0$ and $\ell=2$
In this section, we assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Proposition 5.1 for both $\ell=0$ and $\ell=2$. As before, the condition for $\ell=0$ implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$. For $\ell=2$, we set $g:=T_{2}[\zeta]$, and invoke the corresponding symmetry result from Theorem 6.6, which tells us that
$A_{2}(D)\bigl[g(x)\bigr]$
is odd in the sense of distributions, where $D=\frac{\partial}{\partial x}$ and $A_{2}(D)$ is the differential operator associated to the polynomial
$A_{2}(t)=\frac{t^{2}}{4}-\frac{3t}{2}+3.$
Our goal is again to deduce from this a differential equation satisfied by $\zeta$, and eventually to conclude that $\zeta=0$ under these assumptions. The method follows the same strategy as in the previous section: we express the symmetry condition as an ODE involving $\zeta$, analyze its parity structure, and reduce the problem to a uniqueness statement for solutions of a singular differential equation on $(0,1)$.
Let us set $f:=S_{1}[\zeta]$, so that $g=-S_{2}[f]$. We now differentiate the expression $A_{2}(D)\bigl[g(x)\bigr]$ four times. Since the application of an even number of derivatives preserves parity in the distributional sense, and since the operator $A_{2}(D)$ commutes with differentiation, it follows that
$A_{2}(D)\bigl[f^{(4)}(x)+\tfrac{8}{x}f^{(3)}(x)\bigr]$
is odd, where we have used Lemma 4.1 (vi) with $\ell=2$.
Using again Lemma 4.1 (vi), now with $\ell=1$, and substituting into the above expression, we obtain that the function $A_{2}(D)\bigl[F(x)\bigr]$ is odd in the sense of distributions, where
$F(x):=\zeta^{(4)}(x)+\frac{12}{x}\,\zeta^{(3)}(x)+\frac{24}{x^{2}}\zeta^{\prime\prime}(x)-\frac{24}{x^{3}}\zeta^{\prime}(x).$
As in the cases $\ell=0$ and $\ell=1$, we now define $y(x):=\zeta^{\prime}(x)$. Since $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$, it follows that $y$ is an even function.
Applying the differential operator $4A_{2}(D)=D^{2}-6D+12$ to the function $F(x)$, we define
$G(x):=4A_{2}(D)[F(x)]=\left(\frac{d^{2}}{dx^{2}}-6\frac{d}{dx}+12\right)\left(y^{(3)}+\frac{12}{x}\,y^{\prime\prime}+\frac{24}{x^{2}}y^{\prime}-\frac{24}{x^{3}}y\right).$
We find:
$G(x)=$ $\left(-\frac{288}{x^{5}}-\frac{432}{x^{4}}-\frac{288}{x^{3}}\right)y(x)+\left(\frac{288}{x^{4}}+\frac{432}{x^{3}}+\frac{288}{x^{2}}\right)y^{\prime}(x)$
$+\left(-\frac{96}{x^{3}}-\frac{72}{x^{2}}+\frac{144}{x}\right)y^{\prime\prime}(x)+\left(12-\frac{72}{x}\right)y^{(3)}(x)$
$+\left(-6+\frac{12}{x}\right)y^{(4)}(x)+y^{(5)}(x).$ (8.1)
By the symmetry result, we have that G(x)+G(1x)=0G(x)+G(1-x)=0 in the sense of distributions. This implies that the function $y$ satisfies a linear differential equation of order four, since the fifth derivative $y^{(5)}$ is odd with respect to the midpoint $x=\tfrac{1}{2}$.
More precisely, we obtain the following lemma :
Lemma 8.1.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Lemma 5.1 for $\ell=0$ and $\ell=2$, and set $y(x):=\zeta^{\prime}(x)$. Then $y$ satisfies the differential equation
$\left(\frac{1}{x}+\frac{1}{1-x}-1\right)y^{(4)}(x)-6\left(\frac{1}{x}-\frac{1}{1-x}\right)y^{(3)}(x)$
$\left[12\left(\frac{1}{x}+\frac{1}{1-x}\right)-8\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)-6\left(\frac{1}{x^{2}}+\frac{1}{(1-x)^{2}}\right)\right]y^{\prime\prime}(x)$
$\left[24\left(\frac{1}{x^{2}}-\frac{1}{(1-x)^{2}}\right)+24\left(\frac{1}{x^{4}}-\frac{1}{(1-x)^{4}}\right)+36\left(\frac{1}{x^{3}}-\frac{1}{(1-x)^{3}}\right)\right]y^{\prime}(x)$
$\left[24\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)+24\left(\frac{1}{x^{5}}+\frac{1}{(1-x)^{5}}\right)+36\left(\frac{1}{x^{4}}+\frac{1}{(1-x)^{4}}\right)\right]y(x)=0$
in the sense of distributions on $(0,1)$.
The indicial equation of this differential equation has four solutions, namely $\rho=-2,\,1,\,3,\,4$. As before, we must discard the solution whose asymptotic behavior near $x=0$ is of the form $x^{-2}$.
We shall prove below that there exists a unique even solution in $L^{2}(0,1)$ to this equation. The argument proceeds as follows. Assume, by contradiction, that there exist two linearly independent even solutions. Then, by taking a suitable linear combination, we could find an even solution $y(x)$ such that
$y(0)=y^{\prime}(0)=y^{\prime\prime}(0)=0.$
We now apply the same idea as before, making use of the relation
[…]
which we recompute by performing several integrations by parts. By parity, we first observe that
$\int_{0}^{1}y^{(5)}(x)\,y(x)\,dx=0.$
Moreover, we have the following identities:
$\int_{0}^{1}\left(-6+\frac{12}{x}\right)y^{(4)}(x)\,y(x)\,dx=\int_{0}^{1}\left(-6+\frac{12}{x}\right)y^{\prime\prime}(x)^{2}\,dx\;-\;48\int_{0}^{1}x^{-3}\,y^{\prime}(x)^{2}\,dx\;+\;144\int_{0}^{1}x^{-5}\,y(x)^{2}\,dx,$
$\int_{0}^{1}\left(12-\frac{72}{x}\right)y^{(3)}(x)\,y(x)\,dx=108\int_{0}^{1}x^{-2}\,y^{\prime}(x)^{2}\,dx\;-\;216\int_{0}^{1}x^{-4}\,y(x)^{2}\,dx,$
and
$\int_{0}^{1}P(x)\,y^{\prime\prime}(x)\,y(x)\,dx=-\int_{0}^{1}P(x)\,y^{\prime}(x)^{2}\,dx\;+\;\frac{1}{2}\int_{0}^{1}P^{\prime\prime}(x)\,y(x)^{2}\,dx,$
where
$P(x)\;:=\;-\frac{96}{x^{3}}-\frac{72}{x^{2}}+\frac{144}{x},\qquad P^{\prime\prime}(x)\;=\;-\frac{1152}{x^{5}}-\frac{432}{x^{4}}+\frac{288}{x^{3}}.$
Furthermore,
$\int_{0}^{1}\left(\frac{288}{x^{4}}+\frac{432}{x^{3}}+\frac{288}{x^{2}}\right)y^{\prime}(x)\,y(x)\,dx=\int_{0}^{1}\left(\frac{576}{x^{5}}+\frac{648}{x^{4}}+\frac{288}{x^{3}}\right)y(x)^{2}\,dx.$
Combining all the above identities, we obtain the following compact expression after simplification:
[…]
Now, we make use of the weighted Hardy inequalities in the form
$\int_{0}^{1}\frac{y(x)^{2}}{x^{5}}\,dx\;\leq\;\frac{1}{4}\int_{0}^{1}\frac{y^{\prime}(x)^{2}}{x^{3}}\,dx\,,\qquad\int_{0}^{1}\frac{y(x)^{2}}{x^{4}}\,dx\;\leq\;\frac{4}{9}\int_{0}^{1}\frac{y^{\prime}(x)^{2}}{x^{2}}\,dx\,.$
As a consequence, we deduce that
[…]
We start from (8.2) and observe as above that, for $x\in(0,\frac{1}{2}]$, we have
$f(x)\;:=\;\frac{12}{x^{3}}+\frac{84}{x^{2}}-\frac{144}{x}\geq 96+336-288=144\,.$
Therefore,
$\int_{0}^{1}f(x)\,y^{\prime}(x)^{2}\,dx\;\geq\;144\int_{0}^{\frac{1}{2}}y^{\prime}(x)^{2}dx+\int_{\frac{1}{2}}^{1}f(x)y^{\prime}(x)^{2}dx=\int_{\frac{1}{2}}^{1}(f(x)+144)\,y^{\prime}(x)^{2}dx\,$,
where we have used the symmetry of $y^{\prime}(x)^{2}$ for the last equality. It remains to verify that $f(x)+144$ is positive on $(0,1]$ but its infimum is at $x=1$ and equals $96$.
Since all the integral in (8.2) are positive, we conclude that $y$ must be identically zero. In particular, we have just shown that there exists only one even solution $y(x)\in L^{2}(0,1)$.
The use of a computer is required to determine the unique odd solution $\zeta\in L^{2}(0,1)$ described in Lemma 8.1, (recall that $y(x)=\zeta^{\prime}(x)$). It is explicitly given by
$\zeta(x)\;=\;A\left(4x^{3}-6x^{2}-2x+2-\frac{3(2x-1)}{2\,(x^{2}-x+1)^{2}}+\frac{2(2x-1)}{x^{2}-x+1}\right),$
where $A\in\mathbb{R}$ is an arbitrary constant. Note that the expression $x^{2}-x+1$ is invariant under the transformation $x\mapsto 1-x$.
Unfortunately, this is only a necessary condition, because we applied the operator $A_{2}(D)$ to $T_{2}(\zeta)$ after differentiating four times. We must still check, for this specific $\zeta$, whether $A_{2}(D)\bigl[T_{2}(\zeta)\bigr]$ is indeed an odd function in the sense of distributions.
Recall that, following Rundell and Sacks [23, after Lemma 3.2], the operator $T_{2}$ applied to a function $\zeta$ is defined by
$T_{2}(\zeta)(x)=-\,\zeta(x)-12\,x\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt+24\,x^{3}\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt\,.$
A direct computation yields
$4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr](x)=(288\,x^{3}-432\,x^{2}+144\,x)\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt+(-144\,x+72)\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt$
$\zeta^{\prime\prime}(x)+6\,\zeta^{\prime}(x)-12\,\zeta(x)-\frac{12}{x}\,\zeta^{\prime}(x)+\frac{72}{x}\,\zeta(x)-\frac{48}{x^{2}}\,\zeta(x)\,.$
Evaluating at $x=\tfrac{1}{2}$ (with $A=1$), all integral terms vanish and we obtain, since $\zeta$ is odd,
$4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr]\Big|_{x=\frac{1}{2}}=-60\,\zeta\!\left(\tfrac{1}{2}\right)-18\,\zeta^{\prime}\!\left(\tfrac{1}{2}\right)-\zeta^{\prime\prime}\!\left(\tfrac{1}{2}\right)=-18\,\zeta^{\prime}\!\left(\tfrac{1}{2}\right)=90\,.$
Since $4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr](x)$ is a smooth function, if it were odd with respect to the midpoint $x=\tfrac{1}{2}$, its value at $x=\tfrac{1}{2}$ would necessarily vanish. Therefore, the only possible choice ensuring the oddness of $4\,A_{2}(D)\bigl[T_{2}(\zeta)\bigr]$ in the sense of distributions is to take $A=0$.
We thus obtain a completeness result in the case $(\ell_{1},\ell_{2})=(0,2)$, in full analogy with the case $(\ell_{1},\ell_{2})=(0,1)$:
Theorem 8.2.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Proposition 5.1 for $\ell=0$ and $\ell=2$. Then $\zeta=0$ almost everywhere on $(0,1)$.
In particular, this implies that the spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,2)$ is injective.
9 Conclusion
The techniques developed in this paper rely exclusively on the Kneser–Sommerfeld formula. Even with computer algebra, the analysis of the resulting differential equations is already cumbersome in the cases $(\ell_{1},\ell_{2})=(1,2)$ and $(0,3)$, and larger values currently seem out of reach. This suggests that a more conceptual approach is needed.
We are nevertheless led to the following natural conjecture, originally formulated by Rundell and Sacks: local uniqueness holds near the zero potential as soon as the Dirichlet spectra are known for two distinct values of $\ell$.
Appendix A A Bessel basis of $L^{2}(0,1)$
In [23, Proposition 6.1], Rundell and Sacks established the completeness in $L^{2}(0,1)$ of the family
$\Bigl(1,\ \bigl(\Phi_{0}(n\pi\cdot)\bigr)_{n\geq 1},\ \bigl(\Phi_{2}((n+1)\pi\cdot)\bigr)_{n\geq 1}\Bigr),$ (A.1)
where, for any integer $\ell$,
$\Phi_{\ell}(x):=\tfrac{\pi}{2}\,x\,J_{\ell+\tfrac{1}{2}}(x)^{2}.$
The aim of this appendix is to strengthen their result by proving that the very same family does not merely provide completeness, but in fact constitutes a basis of $L^{2}(0,1)$.
The next lemma will play a key role in the proof of Theorem A.2.
Lemma A.1.
Let $f\in L^{2}(0,1)$. Suppose that both $f$ and $g:=T_{2}^{*}f$ are even with respect to $\tfrac{1}{2}$. Moreover, suppose that $f$ is orthogonal to the constants and to $\cos(2\pi x)$. Then $f=0$.
Proof.
Using Lemma 4.2 successively for $\ell=1$ and $\ell=2$, and recalling that $T_{2}^{*}=-S_{2}^{*}S_{1}^{*}$, we obtain:
$g^{(5)}+\frac{12}{x}\,g^{(4)}+\frac{24}{x^{2}}\,g^{(3)}-\frac{24}{x^{3}}\,g^{\prime\prime}\;=\;-f^{(5)}.$ (A.2)
Projecting this equation onto the subspace of even functions and setting $y=g^{\prime\prime}$, we obtain the following ordinary differential equation:
$y^{\prime\prime}+2\,\frac{1-2x}{x(1-x)}\,y^{\prime}-2\,\frac{3x^{2}-3x+1}{x^{2}(1-x)^{2}}\,y=0.$ (A.3)
The roots of the indicial equation associated with (A.3) are $r=1$ and r=2r=-2. Actually, the general solution of (A.3) on $(0,1)$ can be explicitly written in terms of Ferrers functions (see [18], 14.2) as
$y(x)=g^{\prime\prime}(x)=\frac{C_{1}\,\mathbf{P}^{\,3}_{\nu}(2x-1)+C_{2}\,\mathbf{Q}^{\,3}_{\nu}(2x-1)}{\sqrt{x(1-x)}},\qquad\nu=\frac{\sqrt{33}-1}{2}.$ (A.4)
To enforce the symmetry y(1x)=y(x)y(1-x)=y(x), the coefficients must satisfy (see [18], Section 14.3)
$\frac{C_{2}}{C_{1}}\;=\;\frac{2}{\pi}\,\cot\!\Big(\frac{\pi\nu}{2}\Big).$
An even solution is therefore of the form $y(x)=K\,y_{0}(x)$ for some constant $K\in\mathbb{C}$, where
$y_{0}(x)=\frac{\mathbf{P}^{\,3}_{\nu}(2x-1)\;+\;\tfrac{2}{\pi}\,\cot\!\left(\tfrac{\pi\nu}{2}\right)\,\mathbf{Q}^{\,3}_{\nu}(2x-1)}{\sqrt{x(1-x)}}.$ (A.5)
Using (A.2), we obtain
$y^{(3)}(x)+\frac{12}{x}\,y^{\prime\prime}(x)+\frac{24}{x^{2}}\,y^{\prime}(x)-\frac{24}{x^{3}}\,y(x)\;=\;-\,f^{(5)}(x).$ (A.6)
To simplify notation, set
$z_{0}(x):=y_{0}^{(3)}(x)+\frac{12}{x}\,y_{0}^{\prime\prime}(x)+\frac{24}{x^{2}}\,y_{0}^{\prime}(x)-\frac{24}{x^{3}}\,y_{0}(x).$ (A.7)
Consequently, since $f$ is even with respect to $x=\tfrac{1}{2}$, we may write
$f(x)\;=\;K\,f_{0}(x)\;+\;P(x),$ (A.8)
where
$f_{0}(x)\;=\;-\,\frac{1}{24}\int_{1/2}^{x}(x-t)^{4}\,z_{0}(t)\,dt,$ (A.9)
and
$P(x)\;=\;a\;+\;b\,(x-\tfrac{1}{2})^{2}\;+\;c\,(x-\tfrac{1}{2})^{4}.$
Here $a$, $b$, and $c$ are constants determined by the even Taylor expansion of $f$ at $x=\tfrac{1}{2}$:
$a\;=\;f\!\left(\tfrac{1}{2}\right),\qquad b\;=\;\tfrac{1}{2}\,f^{\prime\prime}\!\left(\tfrac{1}{2}\right),\qquad c\;=\;\tfrac{1}{24}\,f^{(4)}\!\left(\tfrac{1}{2}\right).$ (A.10)
Now, using the definition of $T_{2}^{*}$,
$g(x)=T_{2}^{*}f(x)\;=\;-f(x)-\frac{12}{x^{2}}\int_{0}^{x}tf(t)\,dt\;+\;\frac{24}{x^{4}}\int_{0}^{x}t^{3}f(t)\,dt,$ (A.11)
we readily obtain
$(g+f)^{\prime}(x)$ $=\frac{12}{x}\,f(x)\;+\;\frac{24}{x^{3}}\int_{0}^{x}t\,f(t)\,dt\;-\;\frac{96}{x^{5}}\int_{0}^{x}t^{3}f(t)\,dt,$ (A.12)
$(g+f)^{(3)}(x)$ $=\frac{12}{x}\,f^{\prime\prime}(x)\;-\;\frac{96}{x^{2}}\,f^{\prime}(x)\;+\;\frac{576}{x^{3}}\,f(x)$
$\quad+\;\frac{288}{x^{5}}\int_{0}^{x}t\,f(t)\,dt\;-\;\frac{2880}{x^{7}}\int_{0}^{x}t^{3}f(t)\,dt.$ (A.13)
Consequently, since $f$ and $g$ are even, we obtain at $x=\frac{1}{2}$,
$0\;=\;(g+f)^{\prime}\!\left(\tfrac{1}{2}\right)$ $=24\,a\;+\;192\int_{0}^{1/2}t\,f(t)\,dt\;-\;3072\int_{0}^{1/2}t^{3}f(t)\,dt,$ (A.14)
$0\;=\;(g+f)^{(3)}\!\left(\tfrac{1}{2}\right)$ $=4608\,a\;+\;48\,b\;+\;9216\int_{0}^{1/2}t\,f(t)\,dt\;-\;368640\int_{0}^{1/2}t^{3}f(t)\,dt.$ (A.15)
Substituting $f=Kf_{0}+P$, a straightforward calculation yields:
$b\;+\;\tfrac{1}{7}\,c$ $=K\!\left(-\,960\int_{0}^{1/2}t\,f_{0}(t)\,dt\;+\;15360\int_{0}^{1/2}t^{3}f_{0}(t)\,dt\right),$ (A.16)
$c$ $=K\!\left(26880\int_{0}^{1/2}t\,f_{0}(t)\,dt\;-\;1\,075\,200\int_{0}^{1/2}t^{3}f_{0}(t)\,dt\right).$ (A.17)
From (A.16) and (A.17), we deduce that
$b\;=\;960\,K\!\left(-\,5\int_{0}^{1/2}t\,f_{0}(t)\,dt\;+\;176\int_{0}^{1/2}t^{3}f_{0}(t)\,dt\right).$ (A.18)
Moreover, since $f$ is orthogonal to $\cos(2\pi x)$, we obtain
$b\;+\;\frac{\pi^{2}-6}{2\pi^{2}}\,c\;=\;-\,2\pi^{2}K\int_{0}^{1}f_{0}(x)\,\cos(2\pi x)\,dx.$ (A.19)
A numerical computation with Mathematica yields
[…]
Substituting into (A.16) and (A.17), we deduce that
$b\sim 25.8125\,K,\qquad c\sim-128.293\,K.$
Plugging these values into (A.19), this implies that $K=0$ and hence $b=c=0$. Thus, $f(t)=a$. At last, since $f$ is orthogonal to the constants, $a=0$, which completes the proof of the lemma. ∎
From the preceding lemma, we deduce the following result:
Theorem A.2.
The family
$\left(1,\ \bigl(\Phi_{0}(n\pi\cdot)\bigr)_{n\geq 1},\ \bigl(\Phi_{2}((n+1)\pi\cdot)\bigr)_{n\geq 1}\right)$
forms a basis of $L^{2}(0,1)$.
Proof.
The argument follows the proof of Corollary 5.2 in [23]. Assume that there exist coefficients $\{c_{n}\}_{n\geq 1}$, $\{d_{n}\}_{n\geq 1}$ and $d_{0}\in\mathbb{R}$ such that
$\sum_{n=1}^{\infty}c_{n}\,\Phi_{2}\bigl((n+1)\pi x\bigr)\;+\;d_{0}\;+\;\sum_{n=1}^{\infty}d_{n}\,\Phi_{0}(n\pi x)\;=\;0\quad\text{in }L^{2}(0,1).$ (A.20)
Apply the orthogonal projection $P_{0}$ onto the subspace of functions that are odd with respect to $x=\tfrac{1}{2}$ to both sides. Recalling that $\Phi_{0}(n\pi x)=\sin^{2}(n\pi x)$ is even with respect to $x=\tfrac{1}{2}$, we obtain
$P_{0}\!\left(\sum_{n=1}^{\infty}c_{n}\,\Phi_{2}\bigl((n+1)\pi x\bigr)\right)=0.$ (A.21)
We first show that the family
$\bigl(P_{0}\,\Phi_{2}\bigl((n+1)\pi x\bigr)\bigr)_{n\geq 1}$ (A.22)
is linearly independent. By Lemma 4.3 (ii), this reduces to proving that
$P_{0}\,T_{2}^{*}\!\left(\sum_{n=1}^{\infty}c_{n}\cos\bigl((2n+2)\pi x\bigr)\right)=0\quad\Longrightarrow\quad c_{n}=0\ \text{for all }n.$ (A.23)
Indeed, once $c_{n}=0$ for all $n$, the original relation simplifies to $d_{0}+\sum_{n=1}^{\infty}d_{n}\,\Phi_{0}(n\pi x)=0,$ which forces $d_{n}=0$ for all $n$ (and $d_{0}=0$ as well). This yields the claim.
Define
$f(x)=\sum_{n=1}^{\infty}c_{n}\cos\big((2n+2)\pi x\big).$ (A.24)
The above identity then reads $P_{0}T_{2}^{*}f=0$, that is, both $f$ and $T_{2}^{*}f$ are even with respect to $x=\tfrac{1}{2}$. Moreover, $f$ is orthogonal to the constants and to $\cos(2\pi x)$. By Lemma A.1, it follows that $f\equiv 0$, which proves the result. ∎
Appendix B Proof of Theorem 1.1
In this appendix, we briefly recall the scattering-theoretic framework underlying Theorem 1.1. Throughout this section, we fix the energy $\lambda\in\mathbb{C}$ and consider the radial Schrödinger equation
$u^{\prime\prime}(r)+\left(\frac{\ell(\ell+1)}{r^{2}}+\hat{q}(r)\right)u(r)=\lambda\,u(r),\qquad r>0,$ (B.1)
where $\hat{q}$ denotes the extension of $q\in L^{2}(0,1)$ by zero for $r\geq 1$. [^3] In the remainder of this appendix, we omit the hat notation and use the same notation $q$ for the potential in $L^{2}(0,1)$ and its extension by 0 in $L^{2}(0,+\infty)$.
For each $\nu=\ell+\tfrac{1}{2}$, we denote by $\varphi(r,\nu,\lambda)$ the regular solution of (B.1), characterized by the boundary condition
$\varphi(r,\nu,\lambda)\sim r^{\nu+\frac{1}{2}},\qquad r\to 0.$ (B.2)
By definition, a solution of (B.1) is a $C^{1}$-function whose derivative with respect to $r$ is locally absolutely continuous on $(0,+\infty)$.
For instance, when $q\equiv 0$, the regular solution is explicitly given by
$\varphi_{0}(r,\nu,\lambda)=2^{\nu}\,\lambda^{-\nu/2}\,\Gamma(\nu+1)\,\sqrt{r}\,J_{\nu}(\sqrt{\lambda}\,r),$ (B.3)
where $\sqrt{\lambda}$ is defined by the principal branch. [^4] By convention, if $\lambda=|\lambda|e^{i\theta}$ with $\theta\in(-\pi,\pi]$, then $\sqrt{\lambda}=\sqrt{|\lambda|}\,e^{i\theta/2}$.
The existence and uniqueness of a regular solution are ensured by the following theorem, proved in [12, Lemma 2.2] (see also [13, Lemma B.2]) for locally integrable potentials $q$ on $(0,+\infty)$ under the additional assumption $\int_{0}^{1}r\,|q(r)|\,dr<\infty.$
Theorem B.1.
There exists a unique regular solution $\varphi(r,\nu,\lambda)$ of (B.1) satisfying:
  • (i)
    The map $\lambda\mapsto\varphi(r,\nu,\lambda)$ is an entire function of order $\frac{1}{2}$.
  • (ii)
    There exists $C>0$ such that for all $r>0$ and $\lambda\in\mathbb{C}$,
    $|\varphi(r,\nu,\lambda)|\;\leq\;C\left(\frac{r}{1+|\sqrt{\lambda}|\,r}\right)^{\ell+1}e^{|\Im\sqrt{\lambda}|\,r}.$ (B.4)
  • (iii)
    Moreover, there exists $C>0$ such that for all $r>0$ and $\lambda\in\mathbb{C}$,
    $\bigl|\varphi(r,\nu,\lambda)-\varphi_{0}(r,\nu,\lambda)\bigr|\;\leq\;C\left(\frac{r}{1+|\sqrt{\lambda}|\,r}\right)^{\ell+1}e^{|\Im\sqrt{\lambda}|\,r}\int_{0}^{r}\frac{t\,|q(t)|}{1+|\sqrt{\lambda}|\,t}\,dt.$ (B.5)
The Dirichlet spectrum associated with a fixed angular momentum $\ell$ (or equivalently $\nu=\ell+\tfrac{1}{2}$) is defined as the set of values $\{\lambda_{\ell,n}(q)\}_{n\geq 1}$ such that the regular solution satisfies the boundary condition
$\varphi(1,\nu,\lambda_{\ell,n}(q))=0.$ (B.6)
Equivalently, the Dirichlet spectrum consists of the zeros of the entire function $\lambda\mapsto\varphi(1,\nu,\lambda)$.
It is well known that, for such potentials $q$ and for each fixed angular momentum $\ell\in\mathbb{N}$, the associated Dirichlet spectrum forms a countable sequence of simple, real eigenvalues $\{\lambda_{\ell,n}(q)\}_{n\geq 1}$ diverging to infinity. These facts are classical and can be found in [4, 5].
For instance, when $q\equiv 0$, the Dirichlet eigenvalues are explicitly given by
$\lambda_{\ell,n}(0)=j_{\nu,n}^{2},$ (B.7)
where $j_{\nu,n}$ denotes the $n$-th positive zero of $J_{\nu}(z)$. Recall that the Bessel function $J_{\nu}(z)$ admits the classical Hadamard factorization
$J_{\nu}(z)=\frac{(z/2)^{\nu}}{\Gamma(\nu+1)}\prod_{n=1}^{\infty}\left(1-\frac{z^{2}}{j_{\nu,n}^{2}}\right).$ (B.8)
As a consequence, the free regular solution evaluated at $r=1$ can be written as
$\varphi_{0}(1,\nu,\lambda)=\prod_{n=1}^{\infty}\left(1-\frac{\lambda}{j_{\nu,n}^{2}}\right).$ (B.9)
Similarly, by applying Hadamard’s factorization theorem, $\varphi(1,\nu,\lambda)$ admits the following canonical product representation:
Lemma B.2.
For each fixed $\nu=\ell+\tfrac{1}{2}$ and $q\in L^{2}(0,1)$, the regular solution satisfies
$\varphi(1,\nu,\lambda)=C_{\nu}(q)\ \lambda^{m}\prod_{n=1}^{\infty}\left(1-\frac{\lambda}{\lambda_{\ell,n}(q)}\right),$
where $C_{\nu}(q)\neq 0$ is a constant depending on $\nu$ and $q$, and $m\in\{0,1\}$ denotes the order of the zero at $\lambda=0$.
We next show that equality of the Dirichlet spectra implies equality of the regular solutions evaluated at $r=1$.
Lemma B.3.
Let $\nu=\ell+\tfrac{1}{2}$ and let $q,\tilde{q}\in L^{2}(0,1)$ be two potentials such that their Dirichlet spectra coincide:
$\{\lambda_{\ell,n}(q)\}_{n\geq 1}=\{\tilde{\lambda}_{\ell,n}(\tilde{q})\}_{n\geq 1}.$ (B.10)
Then the corresponding regular solutions satisfy
$\varphi(1,\nu,\lambda;q)=\varphi(1,\nu,\lambda;\tilde{q}),\qquad\forall\,\lambda\in\mathbb{C}.$ (B.11)
Proof.
We denote by a tilde all quantities associated with the potential $\tilde{q}$, and, for brevity, we write $\varphi(1,\nu,\lambda;q)$ simply as $\varphi(1,\nu,\lambda)$. Since the Dirichlet spectra coincide, we have $m=\tilde{m}$. By Theorem B.1 (with $r=1$), for real $\lambda=k^{2}$ and $|k|\to\infty$,
$\bigl|\varphi(1,\nu,\lambda)-\varphi_{0}(1,\nu,\lambda)\bigr|=\mathcal{O}(|k|^{-\ell-2}).$ (B.12)
From the standard asymptotics of Bessel functions, one has as $\lambda\to+\infty$,
$\varphi_{0}(1,\nu,\lambda)\sim c_{\nu}\,k^{-\ell-1}\cos\!\Big(k-\tfrac{\pi\nu}{2}-\tfrac{\pi}{4}\Big),\qquad k=\sqrt{\lambda}\in\mathbb{R}.$ (B.13)
Choosing
$k_{n}=\tfrac{\pi}{2}\Big(\nu+\tfrac{1}{2}\Big)+2\pi n,\qquad\lambda_{n}=k_{n}^{2},$ (B.14)
with $n\in\mathbb{N}$, so that $\cos(\cdot)=1$, we obtain as $n\to+\infty$,
$\varphi(1,\nu,\lambda_{n})\sim\varphi_{0}(1,\nu,\lambda_{n})\sim\tilde{\varphi}(1,\nu,\lambda_{n}).$ (B.15)
[…]
One can also define the Jost solutions $f^{\pm}(r,\nu,\lambda)$ as the unique $C^{1}$ solutions with respect to $r$ of (B.1) satisfying the oscillatory asymptotics
$f^{\pm}(r,\nu,\lambda)\sim e^{\pm i\sqrt{\lambda}\,r},\qquad r\to+\infty.$ (B.16)
The proof of this result is given in [13, Lemma B.4] under the assumptions $\int_{0}^{1}r\,|q(r)|\,dr<\infty$ and $\int_{1}^{+\infty}|q(r)|\,dr<\infty.$ In particular, for $\lambda=1$ (which is the case considered in [6] and [20]), and for such potentials supported in $[0,1]$, one has, for all $r\geq 1$ (see [13, Lemma B.3]),
$f^{+}(r,\nu,1)=e^{i(\nu+\frac{1}{2})\frac{\pi}{2}}\,\sqrt{\frac{\pi r}{2}}\,H_{\nu}^{(1)}(r).$ (B.17)
and
$f^{-}(r,\nu,1)=e^{-i(\nu+\frac{1}{2})\frac{\pi}{2}}\,\sqrt{\frac{\pi r}{2}}\,H_{\nu}^{(2)}(r),$ (B.18)
where $H_{\nu}^{(j)}(r)$ denotes the Hankel function of order $\nu$. It is a fundamental observation that the Jost solutions are entirely independent of the potential for all $r\geq 1$.
The pair of Jost solutions $\{f^{+}(r,\nu,1),\,f^{-}(r,\nu,1)\}$ forms a fundamental system of solutions of (B.1). Hence, at the fixed energy $\lambda=1$, the regular solution can be written as
$\varphi(r,\nu,1)=\alpha(\nu)\,f^{+}(r,\nu,1)+\beta(\nu)\,f^{-}(r,\nu,1),$ (B.19)
where $\alpha(\nu),\,\beta(\nu)\in\mathbb{C}$ are called the Jost functions.
We now recall some basic properties of the Jost functions (see [6] for details). They satisfy the conjugation relation $\overline{\alpha(\nu)}=\beta(\nu)$, and neither $\alpha(\nu)$ nor $\beta(\nu)$ vanishes. This allows us to introduce the Regge interpolation:
$\sigma(\nu)=e^{i\pi(\nu+\frac{1}{2})}\,\frac{\alpha(\nu)}{\beta(\nu)}.$ (B.20)
Note that, following [21], the so-called phase shifts $\delta(\nu)$ are defined as a continuous function of $\nu\in(0,+\infty)$ through the relation
$\sigma(\nu)=e^{2i\,\delta(\nu)}.$
The phase shifts become uniquely determined by imposing the condition $\delta(\nu)\to 0$ as $\nu\to+\infty$, and they can be meromorphically continued to complex values of $\nu$.
Consider now a sequence $(\ell_{k})_{k\geq 1}$ of positive integers satisfying
$\sum_{k}\frac{1}{\ell_{k}}=+\infty,$ (B.21)
that is, the Müntz condition. Then the following uniqueness result holds (see Theorem 2.1 in [20]): if $q,\tilde{q}\in L^{2}(0,1)$ are two potentials whose Regge interpolation functions coincide for $\nu_{k}=\ell_{k}+\tfrac{1}{2}$, then
$q=\tilde{q}\quad\text{a.e. on }(0,1).$ (B.22)
Assume now that $q$ and $\tilde{q}$ share the same Dirichlet spectra $\{\lambda_{\ell_{k},n}(q)\}_{k,n\geq 1}$. Since the Jost solutions $f^{\pm}(r,\nu,1)$ never vanish (see [6, Lemma 4.2]) and are independent of the potential at $r=1$, it follows from (B.19) and Lemma B.3 that the corresponding Regge interpolation functions coincide for $\nu_{k}=\ell_{k}+\tfrac{1}{2}$. This concludes the proof of Theorem 1.1.